# Quasi-alternating surgeries

### v1

@@ Abstract.
We show that all SnapPy census L-space knots admit quasi-alternating surgeries except for the knots $t09847$ and $o9\_30634$ which both do not have any quasi-alternating surgeries. In particular, this finishes Dunfield’s classification of the L-space knots among all SnapPy census knots. In addition, we show that all asymmetric census L-space knots have exactly two quasi-alternating slopes that are consecutive integers. Similar behavior is observed for some of the Baker–Luecke asymmetric L-space knots.
We also classify the quasi-alternating surgeries on torus knots and explore briefly the notion of formal L-space surgeries. This allows us to give examples of asymmetric formal L-spaces.
@@ 1. Introduction
A knot $K$ in $S^{3}$ is called an L-space knot if $K$ admits a Dehn surgery to a Heegaard Floer L-space [OS05]. [^1] A $3$-manifold $M$ is an L-space if its Heegaard Floer homology groups fulfill $\operatorname{rk}(\widehat{\operatorname{HF}}(M))=\chi(\widehat{\operatorname{HF}}(M))$, cf. [KWZ24].
All known examples of L-space knots in $S^{3}$ admit a non-trivial Dehn surgery to an L-space that is the double branched cover of a link in $S^{3}$. Indeed, the first known examples of asymmetric L-space knots in $S^{3}$ were shown to be L-space knots because they admitted a Dehn surgery to the double branched cover of an alternating link [BL20]. It is thus natural that one should wish to study when surgeries on L-space knots are double branched covers over $S^{3}$ and, given such a surgery, what we can say about the branching sets.
@@
[…]
@@
alternating $\implies$ quasi-alternating $\implies$ thin $\implies$ dbc $\cap$ L-space
@@
Here we concentrate on quasi-alternating surgeries, since alternating surgeries are studied extensively in [BKM24a]. Quasi-alternating surgeries are significantly harder to understand and classify than alternating surgeries, which are amenable to study via the use of changemaker lattices [Gre13b, Gre13a, Gre14, Gre15, Gib15, McC15, McC16, McC17a, McC17b].
@@
Our methods work particularly well for asymmetric hyperbolic knots. We classify the quasi-alternating slopes of the asymmetric census knots in §4.2 and explore the quasi-alternating slopes of the asymmetric BL knots in §6. These results, overviewed in §§1.21.3, suggest that potentially any asymmetric hyperbolic L-space knot has exactly two quasi-alternating surgeries and their slopes are consecutive integers, see Question 9.6.
We further examine surgeries yielding formal L-spaces in §8. Introduced in [GL16], formal L-spaces are the natural 3-manifold analogues of quasi-alternating links, see §2.1.3. Indeed, the double branched cover of a quasi-alternating link is always a formal L-space.
@@ 1.1. Existence of quasi-alternating surgeries on the census knots
This article grew from an exploration of the alternating and quasi-alternating slopes of Dunfield’s census knots, the $1267$ hyperbolic knots in $S^{3}$ with complements in the SnapPy census of hyperbolic cusped manifolds that can be triangulated with at most nine ideal tetrahedra [Dun20b]. Almost half of the census knots are L-space knots [Dun20a], and they provide a good testing ground for properties of L-space knots cf. [ABG+23, BKM24a, BKM24b, BBK+24].
More precisely: Among the $1267$ census knots, Dunfield found that $630$ are L-space knots and $635$ are not L-space knots, leaving $2$ with their L-space status undetermined. These are the knots $o9\_30150$ and $o9\_31440$. By demonstrating that they both admit quasi-alternating surgeries, we confirm that these two knots are L-space knots and thereby complete Dunfield’s classification of the L-space knots among the census knots.
@@
The knots $o9\_30150$ and $o9\_31440$ both have a quasi-alternating surgery. Hence they are L-space knots.
@@
Every census L-space knot admits a quasi-alternating surgery, except for the knots $t09847$ and $o9\_30634$ which do not admit any quasi-alternating surgeries.
The full statement of the theorem points out that many of the census L-space knots have exceptional surgeries that are double branched covers of quasi-alternating Montesinos links. Table 4 displays a quasi-alternating surgery for each census L-space knot that does not admit an alternating slope or a quasi-alternating Seifert fibered slope. For the census L-space knots with an alternating surgery, we refer to [BKM24a]. The Seifert fibered surgeries can be found at [Dun20b] or [BKM].
@@ 1.2. Classification of dbc slopes of the asymmetric L-space census knots
For a hyperbolic knot $K$, a slope $r$ is an exceptional slope if $K(r)$ is non-hyperbolic and a symmetry-exceptional slope if $K(r)$ is hyperbolic and the symmetry group of $K(r)$ does not inject into the symmetry group of the complement of $K$. Work of Futer–Purcell–Schleimer [FPS22] enables the computation of a finite list of slopes for a knot $K$ that contains any exceptional or symmetry-exceptional slope, see Corollary 2.6. Since a dbc slope for an asymmetric knot must be an exceptional or symmetry-exceptional slope, this significantly whittles down the search for quasi-alternating slopes of the asymmetric census L-space knots. We determine the symmetry-exceptional slopes for the asymmetric census L-space knots in Theorem 4.1 and present them in Table 1. Ultimately this yields:
@@
Each asymmetric census L-space knot admits exactly two dbc slopes, these two dbc slopes are consecutive integers, and each of these dbc slopes is quasi-alternating but not alternating. These slopes are presented in Tables 4.2 and 3.
In particular, this answers Question 12 from [ABG+23].
For the remaining $623$ census L-space knots that are strongly invertible, the classification of quasi-alternating slopes is not so forthcoming as the set of their slopes that are dbc and L-space is infinite. Indeed, every slope on a strongly invertible knot is a dbc slope, and for an L-space knot $K$, mirrored if needed to make it a positive L-space knot, the set of L-space slopes is given by $[2g(K)-1,\infty)$ [OS11].
@@ 1.3. The Baker–Luecke asymmetric L-space knots
It seems surprising that every asymmetric census knot has exactly two dbc slopes, and these slopes are consecutive integers and both quasi-alternating. One might wonder if this is a general phenomenon, see Question 9.7. Other asymmetric L-space knots were constructed in [BL20]. In [BKM24a] we have shown that the simplest $14$ of these knots have exactly two alternating slopes and they are consecutive integers. Here we generalize these results as follows.
@@
For $6$ of the $14$ simplest asymmetric hyperbolic L-space knots from [BL20], each set of dbc slopes agrees with the set of alternating slopes and consists of two consecutive integers. For the other $8$ knots, each set of dbc slopes contains two consecutive integers which are both alternating slopes and we identify an explicit finite set of further slopes that might be dbc slopes.
@@ 1.4. The classification of quasi-alternating slopes on torus knots
Torus knots punctuate the discrepancy between quasi-alternating slopes and thin slopes. Using results of Issa [Iss18], we classify the set of quasi-alternating slopes of torus knots.
@@
Using Theorem 7.1 with $a=5$, $b=3$, and choosing $m=2$, one finds that a slope $p/q$ for $T_{5,3}$ is quasi-alternating if and only if $p/q>12+1/2$. However, $12+1/2$ is a thin slope: $T_{5,3}(12+1/2)$ is the double branched cover of the knot $K11n50$ which is thin but not quasi-alternating [Gre10]. The classification of thin slopes remains open, see Example 9.3 and Question 9.4.
@@ 1.5. Formal L-space surgeries
It turns out that the set of formal L-space surgeries is always either empty or infinite. This is in contrast to both the quasi-alternating and the alternating slopes which may be non-empty and finite.
@@
In the presence of integral formal L-space surgeries, we have a stronger result.
@@
In general, calculating the set of formal L-space slopes seems challenging. For torus knots, we have the following:
@@
Every non-trivial L-space knot admits L-space surgeries that are not formal L-space surgeries.
@@
Surgery slopes of a knot $K$ in $S^{3}$ are slopes in $\partial\mathcal{N}(K)$, the boundary torus of a tubular neighborhood of $K$. Usually, these are measured with respect to the $S^{3}$-basis $\mu,\lambda$ where $\mu$ denotes the meridian of $K$ and $\lambda$ the Seifert longitude of $K$. (An orientation chosen on $K$ orients $\mu$ and $\lambda$ so that $\mu$ links $K$ positively once and $\lambda$ is coherently parallel to $K$.)
@@
SnapPy uses the geometric basis for cusps of census manifolds (and other hyperbolic manifolds without an alternative basis being implicit). For census knots, tabulated links, and manifolds obtained from link diagrams, SnapPy uses the $S^{3}$-basis. We will follow this convention. For example, the complement of the five crossing twist knot $K5a1$ (in the Hoste–Thistlethwaite notation) is isometric to the census manifold $m015$ while the slope $p/q$ for $K5a1$ is the slope $p/(q-2p)$ for $m015$.
To give surgery descriptions of branching sets for quotients of orientation preserving involutions of manifolds, we use the Hoste–Thistlethwaite tabulation of knots and links as recorded in SnapPy which gives an ordering to the link components. As in SnapPy, surgery information is presented in a list-without-commas of ordered pairs $(p_{i},q_{i})$ indicating $p_{i}/q_{i}$ surgery on the $i$th component (with respect to the $S^{3}$-basis). Here the ordered pair $(0,0)$ is taken to mean that no surgery is performed on the component and it forms part of the branching set. For example, $L11n350(0,0)(-4,1)(0,0)$ describes a link of two components arising from the first and third component of $L11n350$ in the manifold obtained by $-4$ surgery on the second component. When no surgery information is given, the link is just the link in $S^{3}$.
@@ Acknowledgments
We thank Dave Futer, Liam Watson, and Claudius Zibrowius for useful discussions, explanations, and remarks. For the computations performed in this paper, we use and combine KLO (Knot like objects) [Swe], KnotJob [Sch], the Mathematica knot theory package [BNMea], Regina [BBPea], sage [Sag], SnapPy [CDGW], and code developed by Dunfield in [Dun20b, Dun20a]. We thank the creators for making these programs publicly available.
@@
MK is supported by the SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG (Projektnummer 281071066 - TRR 191).
@@ 2.1.1. Tangle exteriors
Here, a tangle exterior $T$ of a knot $K$ in $S^{3}$ is a two-strand tangle in a $3$-ball whose double branched cover is the knot exterior $S^{3}\backslash\mathcal{N}(K)$. Then, via the Montesinos Trick [Mon75], the Dehn surgery $K(r)$ on $K$ by any slope $r$ is diffeomorphic to the double branched cover $\operatorname{DBC}(T(r))$ of the tangle filling $T(r)$ of $T$ by a rational tangle of slope $r$. (The parametrization of slopes for the rational tangle filling of the tangle exterior is inherited from the parametrization of slopes for the Dehn filling of the knot exterior, see for example [Wat17] for more details.)
@@ 2.1.2. Quasi-alternating links
From the definition, it is straightforward to see that any non-split alternating link is quasi-alternating. Furthermore, any non-split quasi-alternating link is Khovanov thin (as well as knot Floer homology thin), implying that its double branched cover is an L-space [OS05, MO08]. On the other hand, there exist quasi-alternating knots that are not alternating [LM], thin knots that are not quasi-alternating [Gre10], cf. [GW13], and thick knots whose double branched covers are L-spaces, see for example [Iss18].

### v2

@@ Abstract.
We show that all SnapPy census L-space knots admit quasi-alternating surgeries except for the knots $t09847$ and o9_30634o9\_30634, neither of which have any quasi-alternating surgeries. In particular, this finishes Dunfield’s classification of the L-space knots among all SnapPy census knots. In addition, we show that all asymmetric census L-space knots have exactly two quasi-alternating slopes and that these are consecutive integers. Similar behavior is observed for some of the Baker–Luecke asymmetric L-space knots.
We also classify the quasi-alternating surgeries on torus knots and show that the set of formal L-space slopes is either empty or infinite This allows us to give examples of asymmetric formal L-spaces.
@@ 1. Introduction
A knot $K$ in $S^{3}$ is called an L-space knot if $K$ admits a Dehn surgery to a Heegaard Floer L-space [OS05]. [^1] A $3$-manifold $M$ is an L-space if its Heegaard Floer homology groups fulfill $\operatorname{rk}(\widehat{\operatorname{HF}}(M))=\chi(\widehat{\operatorname{HF}}(M))$, cf. [kotelskiy2022thin].
All known examples of L-space knots in $S^{3}$ admit a non-trivial Dehn surgery to an L-space that is the double branched cover of a link in $S^{3}$. Indeed, the first known examples of asymmetric [^2] Recall that a $3$-manifold is called asymmetric if its mapping class group is trivial. A knot is called asymmetric, if its exterior is asymmetric. L-space knots in $S^{3}$ were shown to be L-space knots because they admitted a Dehn surgery to the double branched cover of an alternating link [BL17]. It is thus natural that one should wish to study when surgeries on L-space knots are a double branched cover of a link in $S^{3}$ and, given such a surgery, what we can say about the branching sets.
@@
{non-split alternating} $\subsetneq$ {quasi-alternating} $\subsetneq$ {Khovanov thin}
@@
alternating [^3] The double branched cover of a split link is reducible with an $S^{1}\times S^{2}$ summand. However, Property R [Ga87] implies that only the unknot has such a surgery among knots in $S^{3}$. $\implies$ quasi-alternating $\implies$ thin $\implies$ dbc $\cap$ L-space [^4] It follows from [OS05] that the double branched cover of a thin knot is an L-space, see for example Theorem 4.5 in [Marengon_thesis].
@@
Here we concentrate on quasi-alternating surgeries, since alternating surgeries are studied extensively in [BKM_alt]. Quasi-alternating surgeries are significantly harder to understand and classify than alternating surgeries, which are amenable to study via the use of changemaker lattices [Gr13, GreeneLRP, Gr14, Greene2015genusbounds, Gi15, Mc15, Mc16, McCoyAltUnknotting, Mc17].
@@
Our methods work particularly well for asymmetric hyperbolic knots. We classify the quasi-alternating slopes of the asymmetric census knots in §4.2 and explore the quasi-alternating slopes of the asymmetric Baker–Luecke knots in §6. These results, overviewed in §§1.21.3, suggest that potentially any asymmetric hyperbolic L-space knot has exactly two quasi-alternating surgeries and their slopes are consecutive integers, see Question 9.7.
We further examine surgeries yielding formal L-spaces in §8. Introduced in [Greene2016Strong], formal L-spaces are the natural 3-manifold analogues of quasi-alternating links, see §2.2.2. Indeed, the double branched cover of a quasi-alternating link is always a formal L-space.
@@ 1.1. Existence of quasi-alternating surgeries on the census knots
This article grew from an exploration of the alternating and quasi-alternating slopes of Dunfield’s census knots, the $1267$ hyperbolic knots in $S^{3}$ with complements in the SnapPy census of hyperbolic cusped manifolds that can be triangulated with at most nine ideal tetrahedra [Du18]. Almost half of the census knots are L-space knots [Du19], and they provide a good testing ground for properties of L-space knots cf. [ABG+19, BKM_alt, BKM_thin, CensusKnotInvariants].
More precisely: Among the $1267$ census knots, Dunfield found that $630$ are L-space knots and $635$ are not L-space knots, leaving $2$ with their L-space status undetermined. These are the knots o9_30150o9\_30150 and o9_31440o9\_31440. By demonstrating that they both admit quasi-alternating surgeries, we confirm that these two knots are L-space knots and thereby complete Dunfield’s classification of the L-space knots among the census knots.
@@
The knots o9_30150o9\_30150 and o9_31440o9\_31440 both have a quasi-alternating surgery. Hence they are L-space knots.
@@
Every census L-space knot admits a quasi-alternating surgery, except for the knots $t09847$ and o9_30634o9\_30634 which do not admit any quasi-alternating surgeries.
The full statement of the theorem points out that many of the census L-space knots have surgeries to Seifert fibered spaces that are double branched covers of quasi-alternating Montesinos links. Table 5 displays a quasi-alternating surgery for each census L-space knot that does not admit an alternating slope or a quasi-alternating Seifert fibered slope. For the census L-space knots with an alternating surgery, we refer to [BKM_alt]. The Seifert fibered surgeries can be found at [Du18] or [BKM].
@@ 1.2. Classification of dbc slopes of the asymmetric L-space census knots
For a hyperbolic knot $K$, a slope $r$ is an exceptional slope if $K(r)$ is non-hyperbolic and a symmetry-exceptional slope if $K(r)$ is hyperbolic and the symmetry group of $K(r)$ does not inject into the symmetry group of the complement of $K$. Work of Futer–Purcell–Schleimer [FPS19] enables the computation of a finite list of slopes for a knot $K$ that contains any exceptional or symmetry-exceptional slope, see Corollary 2.7. Since a dbc slope for an asymmetric knot must be an exceptional or symmetry-exceptional slope, this significantly whittles down the search for quasi-alternating slopes of the asymmetric census L-space knots. We determine the symmetry-exceptional slopes for the asymmetric census L-space knots in Theorem 4.1 and present them in Table 2. Ultimately this yields:
@@
Each asymmetric census L-space knot admits exactly two dbc slopes, these two dbc slopes are consecutive integers, and each of these dbc slopes is quasi-alternating but not alternating. These slopes are presented in Tables 3 and 4.
In particular, this answers Question 12 from [ABG+19].
For the remaining $623$ census L-space knots that are strongly invertible, the classification of quasi-alternating slopes is not so forthcoming as the set of their slopes that are dbc and L-space is infinite. Indeed, every slope on a strongly invertible knot is a dbc slope, and for an L-space knot $K$, mirrored if needed to make it a positive L-space knot, the set of L-space slopes is given by $2g(K)-1,\infty) [$genus_L_spaces].
@@ 1.3. The Baker–Luecke asymmetric L-space knots
It seems surprising that every asymmetric census knot has exactly two dbc slopes, and these slopes are consecutive integers and both quasi-alternating. One might wonder if this is a general phenomenon, see Question 9.7. Other asymmetric L-space knots were constructed in [BL17]. In [BKM_alt] we have shown that the simplest $14$ of these knots have exactly two alternating slopes and they are consecutive integers. Here we generalize these results as follows.
@@
For $6$ of the $14$ simplest asymmetric hyperbolic L-space knots from [BL17], the set of dbc slopes agrees with the set of alternating slopes and consists of two consecutive integers. For the other $8$ knots, the set of dbc slopes contains two consecutive integers which are both alternating slopes and we identify an explicit finite set of further slopes that might be dbc slopes.
@@ 1.4. The classification of quasi-alternating slopes on torus knots
Torus knots punctuate the discrepancy between quasi-alternating slopes and thin slopes. Using results of Issa [Is18], we classify the set of quasi-alternating slopes of torus knots.
@@
Using Theorem 7.1 with $a=5$, $b=3$, and choosing $m=2$, one finds that a slope $p/q$ for $T_{5,3}$ is quasi-alternating if and only if $p/q>12+1/2$. However, $12+1/2$ is a thin slope: $T_{5,3}(12+1/2)$ is the double branched cover of the knot $K11n50$ which is thin but not quasi-alternating [Greene_thin_QA]. The classification of thin slopes remains open, see Example 9.3 and Question 9.4.
@@ 1.5. Formal L-space surgeries
Since the double branched cover of a quasi-alternating link is an example of a formal L-space (see Section 2.2.2 for the definition), it is natural to compare the formal L-spaces surgeries to the quasi-alternating surgeries. It turns out that the set of formal L-space slopes is always either empty or infinite. This is in contrast to both the quasi-alternating and the alternating slopes which may be non-empty and finite.
@@
Remark 1.2
In the proof of Corollary 8.2 we will see that there exist infinitely many formal asymmetric L-spaces that arise by surgery on a knot. In particular, there exist formal L-space slopes that are not thin slopes. On the other hand, the slope $12+1/2$ is thin for $T_{5,3}$ as discussed in Example 1.1, but in [Greene_thin_QA] it is shown that this slope is not a formal L-space slope. Thus thin slopes and formal L-space slopes are in general different.
In the presence of integral formal L-space slopes, we have a stronger result.
@@
In general, calculating the full set of formal L-space slopes seems challenging. However, for torus knots, it turns out to coincide with the set of quasi-alternating surgeries.
@@
Recall from [OS05b] that if $K$ is an L-space knot, then $K(r)$ is an L-space if and only if $r\geq 2g(K)-1$. Thus the following result implies that every non-trivial L-space knot admits L-space slopes that are not formal L-space slopes. In particular, there is no non-trivial L-space knot for which every L-space surgery is a quasi-alternating surgery.
@@
Surgery slopes of a knot $K$ in $S^{3}$ are slopes in $\partial\mathcal{N}(K)$, the boundary torus of a tubular neighborhood of $K$. Usually, these are measured with respect to the $S^{3}$-basis $\mu,\lambda$ where $\mu$ denotes the meridian of $K$ and $\lambda$ the Seifert longitude of $K$. (An orientation chosen on $K$ orients $\mu$ and $\lambda$ so that $\mu$ links $K$ positively once and the pair $(\mu,\lambda)$ yields the positive orientation of $\partial\mathcal{N}(K)$.)
@@
SnapPy uses the geometric basis for cusps of census manifolds (and other hyperbolic manifolds without an alternative basis being implicit). For census knots, tabulated links, and manifolds obtained from link diagrams, SnapPy uses the $S^{3}$-basis. We will follow this convention and use different slope indices according to how the knot/link is described. For example, the complement of the five-crossing twist knot $K5a1$ (in the Hoste–Thistlethwaite notation) is isometric to the census manifold $m015$ while the slope $p/q$ for $K5a1$ is the slope p/(q2p)p/(q-2p) for $m015$.
To give surgery descriptions of quotients of orientation preserving involutions of manifolds, we use the Hoste–Thistlethwaite–Weeks tabulation of knots and links [HTW] (in short: the HTW table) as recorded in SnapPy which gives an ordering to the link components. As in SnapPy, surgery information is presented in a list-without-commas of ordered pairs $(p_{i},q_{i})$ indicating $p_{i}/q_{i}$ surgery on the $i$th component (with respect to the $S^{3}$-basis). Here the ordered pair $(0,0)$ is taken to mean that no surgery is performed on the component and it forms part of the branching set. For example, L11n350(0,0)(4,1)(0,0)L11n350(0,0)(-4,1)(0,0) describes a link of two components arising from the first and third component of $L11n350$ in the manifold obtained by 4-4 surgery on the second component. When no surgery information is given, the link is just the link in $S^{3}$.
@@ Acknowledgments
We thank Dave Futer, Liam Watson, and Claudius Zibrowius for useful discussions, explanations, and remarks. For the computations performed in this paper, we use and combine KLO (Knot like objects) [Sw], KnotJob [KnotJob], the Mathematica knot theory package [knotatlas], Regina [BBP+], sage [sagemath], SnapPy [CDGW], and code developed by Dunfield in [Du18, Du19]. We thank the creators for making these programs publicly available. We are also grateful to the three anonymous referees for careful reading, useful suggestions, and comments.
@@
MK was supported by the SFB/TRR 191 Symplectic Structures in Geometry, Algebra and Dynamics, funded by the DFG, German Research Foundation (Projektnummer 281071066 - TRR 191), by the DFG (Project: 561898308); by a Ramón y Cajal grant (RYC2023-043251-I) and by the project PID2024-157173NB-I00 funded by MCIN/AEI/10.13039/501100011033, ESF+ and FEDER, EU; and by a VII Plan Propio de Investigación y Transferencia (SOL2025-36103) of the University of Sevilla.
@@ 2.1.1. Tangle exteriors
Here, a tangle exterior $T$ of a knot $K$ in $S^{3}$ is a two-strand tangle in a $3$-ball whose double branched cover is the knot exterior $S^{3}\backslash\mathcal{N}(K)$. Then, via the Montesinos Trick [Montesinos_trick], the Dehn surgery $K(r)$ on $K$ by any slope $r$ is diffeomorphic to the double branched cover $\operatorname{DBC}(T(r))$ of the tangle filling $T(r)$ of $T$ by a rational tangle of slope $r$. (The parametrization of slopes for the rational tangle filling of the tangle exterior is inherited from the parametrization of slopes for the Dehn filling of the knot exterior, see for example [Watson_KH_sym2017] for more details.)
2.2. Thin links
A link $L$ is called thin, if the Khovanov homology $Kh^{h,q}(L)$ (with $\mathbb{Z}_{2}$-coefficients) of $L$ is supported in two diagonal gradings $\delta=q-2h$. While the Khovanov homology of a link depends on the orientation of its components, the thinness status of a link is independent of the orientations of the link components. This follows, for example, from Proposition 28 in [Khovanov_homology]. A link that is not thin is called thick.
@@ 2.2.1. Quasi-alternating links
From the definition, it is straightforward to see that any non-split alternating link is quasi-alternating. Furthermore, any non-split quasi-alternating link is Khovanov thin (as well as knot Floer homology thin), implying that its double branched cover is an L-space [OS05, QA_thin]. On the other hand, there exist quasi-alternating knots that are not alternating [KnotInfo], thin knots that are not quasi-alternating [Greene_thin_QA], cf. [GW13], and thick knots whose double branched cover is an L-space, see for example [Is18]. [^5] As pointed out by a referee, a simple example of a thick know whose double branched cover is an L-space is given by the torus knot $T_{3,5}$.